Can someone help please? thanks!
Simplify each expression :
64^7/6
thankyou.

Answer by jim_thompson5910(35256)

64^7/6 = (64^1/6)^7

64^7/6 = (6th root of 64)^7

64^7/6 = 2^7

64^7/6 = 128